### v1

@@ 4.3.1 Optimal decision tree algorithms for size-constrained tree
Proof.
The result is a simple consequence of distributivity
@@
where function $\mathit{ins}$ insert $r$ into all possible positions for each partial permutations in $\mathit{perms}\left(\mathit{rs}\right)$ ( see subsection II.1.2.4 in He (2025) for details). Comparing (26) with (25), we can define $\mathit{sodt}$ using $\mathit{perms}$ through the following equivalence
@@
$E\left(\mathit{mapD}{}_{\cap_{r^{+}}}\left(u\right)\right)\leq E\left(\mathit{mapD}{}_{\cap_{r^{\prime+}}}\left(u^{\prime}\right)\right)\wedge E\left(\mathit{mapD}{}_{\cap_{r^{-}}}\left(v\right)\right)\leq E\left(\mathit{mapD}{}_{\cap_{r^{\prime-}}}\left(v^{\prime}\right)\right)\Longrightarrow E\left(DN\left(u,r,v\right)\right)\leq E\left(DN\left(u^{\prime},r^{\prime},v^{\prime}\right)\right)$ (32)
@@
Astute readers may have noticed that the recursive structure of $E$ aligns with that of $\mathit{genDTs}_{\text{rec}}$. We now show that this alignment implies a dynamic programming (DP) solution. By leveraging the monotonicity property 2, we can fuse the minimizer $\mathit{min}_{E}$ with $\mathit{genDTs}_{\text{rec}}$ in the simplified decision tree problem defined in 28. This yields a DP solution $\mathit{sodt}_{\text{rec}}$ for the problem $\mathit{min}_{E}\circ\mathit{genDTs}_{\text{rec}}$ 28. Thus providing an efficient algorithm for solving the optimal decision tree problem by substituting the DP solution $\mathit{sodt}_{\text{rec}}$ into $\mathit{odt}_{\text{rec}}$.
Theorem 3.
@@
$\mathit{sodt}_{\text{rec}}\left(\mathit{xs}\right)\subseteq\mathit{min}_{E}\circ\mathit{genDTs}_{\text{rec}}\left(\mathit{xs}\right),$
@@
Proof.
Due to the monotonicity of the problem, we can now derive the program by following equational reasoning
@@ 4.3.2 Optimal decision tree algorithm for depth-constrained tree
A similar reasoning applies to the ODT problem with depth constraints, commonly used in axis-parallel and binary-feature decision trees. Brița et al. (2025) present an algorithm for this problem; however, instead of formally deriving it, they present it directly. While their approach can be seen as a special case of the algorithm derived here, a formal derivation is important for understanding the underlying principles and for establishing a rigorous foundation.
The following theorem explains why the algorithm of Brița et al. (2025) is correct and how it can be systematically derived from (24).
Theorem 4.
@@
Proof.
The base cases ($d=1$ and $d=0$) follow directly from the definition of an optimal decision tree:
@@
$\mathit{min}_{E}\left[\mathit{DN}\left(u,r,v\right)\mid u\leftarrow\mathit{genDTDs}_{\text{depth}}\left(d-1,\mathit{xs}^{+}\right),v\leftarrow\mathit{genDTDs}_{\text{depth}}\left(d-1,\mathit{xs}^{-}\right),r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\right]$
@@
$\left[\mathit{DN}\left(\mathit{min}_{E}\left(\mathit{genDTDs}_{\text{depth}}\left(d-1,\mathit{xs}^{+}\right)\right),r,\mathit{min}_{E}\left(\mathit{genDTDs}_{\text{depth}}\left(d-1,\mathit{xs}^{-}\right)\right)\right)\mid r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\right]$
@@
This proved that $\mathit{odt}_{\text{depth}}\left(d,\mathit{xs}\right)$ can be computed recursively using optimal subtrees for $\mathit{xs}^{+}$ and $\mathit{xs}^{-}$ of depth d1d-1. ∎
Interestingly, for AODT and ODT over binary-feature data, Brița et al. (2025) and Demirović et al. (2022) observed that $\mathit{odt}_{\text{depth}}\left(2,\mathit{xs}\right)$ can be solved more efficiently than the classical recursive method. Expressing $\mathit{odt}_{\text{depth}}\left(2,\mathit{xs}\right)$ in a more compact form and substituting it into the recursion 34, may help convey their algorithmic insights more clearly.
@@ 4.3.3 Non-existence of monotonicity
Proof.
It suffices to show that the following implication does not hold:
@@
where $\mathit{insR}:\left[\mathcal{R}\right]\to\left[\mathcal{R}\right]$ is the relational version of the $\mathit{ins}$ operation He (2025), which inserts $r$ to a random position in a existing permutation of rules $\mathit{rs}$. The implication fails because the newly inserted rule $r$ can become the root of the resulting tree, which can change the error non-monotonically.
Therefore, dynamic programming or greedy algorithms do not exist for $\mathit{sodt}_{\text{kperms}}$ and $\mathit{sodt}_{\text{vec}}$, when the objective function is defined according to (30). However, this does not imply that DP or greedy solutions are impossible for objectives defined in other form. For instance, if the objective function is incrementally definable—such as the tree size—then a greedy solution exist. ∎
@@ 4.3.4 The use of memoization or caching
Caching provides a practical alternative when a bottom-up recursion or iterative DP is either unavailable or too complex to implement efficiently. This approach has been widely adopted in ODT algorithms over binary feature data. For instance, Aglin et al. (2021) employ caches over itemsets (collections of Boolean features), while Demirović et al. (2022) note a trade-off between branch and dataset caching, as introduced by Nijssen and Fromont (2007).
The correctness of these caching techniques in ODT studies remains questionable. The derivations of $\mathit{odt}_{\text{depth}}$ and $\mathit{odt}_{\text{size}}$ depend on Lemma 2, which holds only when the roots of two trees, $r$ and $r^{\prime}$ are equivalent. Therefore, correct use of memoization requires storing both the root and its corresponding subtrees, because optimality of subtrees $u$ and $v$ over subregions $r^{\pm}$ implies optimality of the full tree $DN\left(u,r,v\right)$. Thus, no trade-off exists between branch and dataset caching: both subtrees (depth- or size-constrained) and datasets must be stored. In this context, the caching strategy in Brița et al. (2025) appears correct, as their ConTree algorithm reuses cached solutions defined both by sub-dataset and subtree depth.
@@
$O\left(\left|\mathcal{S}_{\mathit{\text{split}}}\left(\mathinner{xs}\right)\right|\times\sum_{k=0}^{K-1}\left|\mathcal{S}_{\text{size}}\left(k,\mathinner{xs}\right)\right|\right)$
@@
$O\left(\left|\mathcal{S}_{\mathit{\text{split}}}\left(\mathinner{xs}\right)\right|\times\sum_{i=0}^{d-1}\left|\mathcal{S}_{\text{depth}}\left(i,\mathinner{xs}\right)\right|\right)$
@@
This results in formidable combinatorial complexity for most ODT problems. For example, in the AODT problem, the number of possible rules is $\mathcal{S}_{\mathcal{H}^{0}}=O\left(N\times D\right)$. For hyperplane ODT problems, the number of possible splits grows to $\mathcal{S}_{\mathcal{H}}=O\left(N^{{}^{D}}\right)$ (see Section 5). Consequently, the caching approach in Brița et al. (2025), which stores only a few thousand optimal sub-trees, has a very low probability of cache hits. Our tests on multiple data lists confirm this: in extreme cases, their algorithm examined millions of decision trees without achieving a single cache hit. This suggests that the efficiency claimed for their ConTree algorithm due to caching is overstated; most gains likely arise from factors other than caching.

### v2

@@ 4.3.1 Optimal decision tree algorithms for size-constrained tree
Proof The result is a simple consequence of distributivity
@@
where function $\mathit{ins}$ insert $r$ into all possible positions for each partial permutations in $\mathit{perms}\left(\mathit{rs}\right)$ ( see subsection II.1.2.4 in He (2025b) for details). Comparing (27) with (26), we can define $\mathit{sodt}$ using $\mathit{perms}$ through the following equivalence
@@
$E\left(\mathit{mapD}_{\cap_{r^{+}}}(u)\right)\leq$ $E\left(\mathit{mapD}_{\cap_{r^{\prime+}}}(u^{\prime})\right)\wedge E\left(\mathit{mapD}_{\cap_{r^{-}}}(v)\right)\leq E\left(\mathit{mapD}_{\cap_{r^{\prime-}}}(v^{\prime})\right)$ (34)
$\quad\Longrightarrow E\left(DN(u,r,v)\right)\leq E\left(DN(u^{\prime},r^{\prime},v^{\prime})\right)$
@@
Astute readers may have noticed that the recursive structure of $E$ aligns with that of $\mathit{genDTs}_{\text{rec}}$. We now show that this alignment implies a dynamic programming (DP) solution. By leveraging the monotonicity property 6, we can fuse the minimizer $\mathit{min}_{E}$ with $\mathit{genDTs}_{\text{rec}}$ in the simplified decision tree problem defined in 30. This yields a DP solution $\mathit{sodt}_{\text{rec}}$ for the problem $\mathit{min}_{E}\circ\mathit{genDTs}_{\text{rec}}$ 30. Thus providing an efficient algorithm for solving the optimal decision tree problem by substituting the DP solution $\mathit{sodt}_{\text{rec}}$ into $\mathit{odt}_{\text{rec}}$.
Theorem 7
@@
$\mathit{sodt}_{\text{rec}}\left(\mathit{xs}\right)\subseteq\mathit{min}_{E}\circ\mathit{genDTs}_{\text{rec}}\left(\mathit{xs}\right),$ (36)
@@
Proof Due to the monotonicity of the problem, we can now derive the program by following equational reasoning
@@ 4.3.2 Optimal decision tree algorithm for depth-constrained tree
A similar reasoning applies to the ODT problem with depth constraints, commonly used in axis-parallel and binary-feature decision trees. Brita et al. (2025) present an algorithm for this problem; however, instead of formally deriving it, they present it directly. While their approach can be seen as a special case of the algorithm derived here, a formal derivation is important for understanding the underlying principles and for establishing a rigorous foundation.
The following theorem explains why the algorithm of Brita et al. (2025) is correct and how it can be systematically derived from (25).
Theorem 8
@@
Proof The base cases ($d=1$ and $d=0$) follow directly from the definition of an optimal decision tree:
@@
$\mathit{min}_{E}\big[\mathit{DN}\left(u,r,v\right)\mid u\leftarrow\mathit{genDTDs}_{\text{depth}}\left(d-1,\mathit{xs}^{+}\right),$
$\quad\quad v\leftarrow\mathit{genDTDs}_{\text{depth}}\left(d-1,\mathit{xs}^{-}\right),r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\big]$
@@
$\big[\mathit{DN}\big(\mathit{min}_{E}\left(\mathit{genDTDs}_{\text{depth}}\left(d-1,\mathit{xs}^{+}\right)\right),$
$\quad\quad r,\mathit{min}_{E}\left(\mathit{genDTDs}_{\text{depth}}\left(d-1,\mathit{xs}^{-}\right)\right)\big)\mid r\leftarrow\mathit{gen_{splits}}\left(\mathit{xs}\right)\big]$
@@
This proved that $\mathit{odt}_{\text{depth}}\left(d,\mathit{xs}\right)$ can be computed recursively using optimal subtrees for $\mathit{xs}^{+}$ and $\mathit{xs}^{-}$ of depth d1d-1.
Interestingly, for AODT and ODT over binary-feature data, Brita et al. (2025) and Demirovic et al. (2022) observed that $\mathit{odt}_{\text{depth}}\left(2,\mathit{xs}\right)$ can be solved more efficiently than the classical recursive method. Expressing $\mathit{odt}_{\text{depth}}\left(2,\mathit{xs}\right)$ in a more compact form and substituting it into the recursion 37, may help convey their algorithmic insights more clearly.
@@ 4.3.3 Non-existence of monotonicity
Proof It suffices to show that the following implication does not hold:
@@
where $\mathit{insR}:\left[\mathcal{R}\right]\to\left[\mathcal{R}\right]$ is the relational version of the $\mathit{ins}$ operation He (2025b), which inserts $r$ to a random position in a existing permutation of rules $\mathit{rs}$. The implication fails because the newly inserted rule $r$ can become the root of the resulting tree, which can change the error non-monotonically.
Therefore, dynamic programming or greedy algorithms do not exist for $\mathit{sodt}_{\text{kperms}}$ and $\mathit{sodt}_{\text{vec}}$, when the objective function is defined according to (32). However, this does not imply that DP or greedy solutions are impossible for objectives defined in other form. For instance, if the objective function is incrementally definable—such as the tree size—then a greedy solution exist.
@@ 4.3.4 The use of memoization or caching
Caching provides a practical alternative when a bottom-up recursion or iterative DP is either unavailable or too complex to implement efficiently. This approach has been widely adopted in ODT algorithms over binary feature data. For instance, Aglin et al. (2021) employ caches over itemsets (collections of Boolean features), while Demirovic et al. (2022) note a trade-off between branch and dataset caching, as introduced by Nijssen and Fromont (2007).
The correctness of these caching techniques in ODT studies remains questionable. The derivations of $\mathit{odt}_{\text{depth}}$ and $\mathit{odt}_{\text{size}}$ depend on Lemma 6, which holds only when the roots of two trees, $r$ and $r^{\prime}$ are equivalent. Therefore, correct use of memoization requires storing both the root and its corresponding subtrees, because optimality of subtrees $u$ and $v$ over subregions $r^{\pm}$ implies optimality of the full tree $DN\left(u,r,v\right)$. Thus, no trade-off exists between branch and dataset caching: both subtrees (depth- or size-constrained) and datasets must be stored. In this context, the caching strategy in Brita et al. (2025) appears correct, as their ConTree algorithm reuses cached solutions defined both by sub-dataset and subtree depth.
@@
$O\left(\left|\mathcal{S}_{\mathit{\text{split}}}\left(\mathinner{xs}\right)\right|\times\sum_{k=0}^{K-1}\left|\mathcal{S}_{\text{size}}\left(k,\mathinner{xs}\right)\right|\right)$ (38)
@@
$O\left(\left|\mathcal{S}_{\mathit{\text{split}}}\left(\mathinner{xs}\right)\right|\times\sum_{i=0}^{d-1}\left|\mathcal{S}_{\text{depth}}\left(i,\mathinner{xs}\right)\right|\right)$ (39)
@@
This results in formidable combinatorial complexity for most ODT problems. For example, in the AODT problem, the number of possible rules is $\mathcal{S}_{\mathcal{H}^{0}}=O\left(N\times D\right)$. For hyperplane ODT problems, the number of possible splits grows to $\mathcal{S}_{\mathcal{H}}=O\left(N^{{}^{D}}\right)$ (see Section 5). Consequently, the caching approach in Brita et al. (2025), which stores only a few thousand optimal sub-trees, has a very low probability of cache hits. Our tests on multiple data lists confirm this: in extreme cases, their algorithm examined millions of decision trees without achieving a single cache hit. This suggests that the efficiency claimed for their ConTree algorithm due to caching is overstated; most gains likely arise from factors other than caching.
4.4 Extension to decision tree problem over binary feature data (non-proper decision tree)
4.4.1 Definition of the search space
A key advantage of adapting a general axiomatic framework for defining decision trees is that it provides an unambiguous method for classifying decision tree problems. The optimal decision tree problem over binary feature data (ODT-BF) aims to construct a tree that encodes a mapping over binary feature datasets, i.e., $t:\left\{-1,1\right\}^{D}\to L$, where $L$ denotes the set of labels.
The ODT-BF problem is perhaps the most extensively studied case of applying combinatorial methods Hu et al. (2019); Lin et al. (2020); Zhang et al. (2023); Demirovic et al. (2022); Aglin et al. (2021, 2020); Nijssen and Fromont (2007, 2010). The tree structure of ODT-BF is defined in the same way as proper decision trees; in other words, ODT-BF also satisfies the structural constraints of decision trees (Axioms 1–3) and is therefore a valid decision tree problem. However, this problem has much simpler combinatorics, and indeed He (2025b) showed that its complexity is independent of the input data size, and empirical results from Hu et al. (2019) further indicate that, when the feature dimension is fixed, the algorithm scales linearly with data size.
Characterization of splitting rules
Unlike decision tree problems over continuous datasets, where splitting rules can often be interpreted as geometric partitions—such as a hyperplane dividing the space into two regions—ODT-BF cannot naturally be viewed this way. Some researchers classify ODT-BF within the class of axis-parallel hyperplanes, but this characterization is misleading, as it obscures the geometric meaning of axis-parallel splits. Binary feature data correspond to vertices of a hypercube, and splits cannot be defined by midpoints between vertices, since the data are discrete rather than continuous.
PA more natural interpretation is to view each splitting rule in ODT-BF as a logical question. Specifically, given a data point $x\in\mathbb{R}^{D}$, each splitting rule in ODT-BF problem corresponds to determining whether the $i$-th feature of a data point $x\in\mathbb{R}^{D}$ is 0 or 1—equivalently, whether $x$ contains feature $i$?
An immediate consequence of this characterization is that the number of splitting rules is independent of the dataset size, as a hypercube in $\mathbb{R}^{D}$ has at most $2^{D}$ vertexes. Indeed, in the ODT-BF problem, many data points are duplicates and occupy the same vertices of the cube, and there are many bounding techniques are developed based on this observation (Lin et al., 2020; Zhang et al., 2023; Hu et al., 2019).
Ancestral constraints and splitting function
Having characterized the splitting rules in ODT-BF as simple logical questions, we now analyze the ancestral constraints between pairs of rules. Suppose a rule $r_{i}$ is an ancestor of another rule $r_{j}$. The question of whether $x$ contains feature $i$ is independent of whether it contains feature $j$. This means that fixing an ancestor rule $r_{i}$ does not influence whether $r_{j}$ should go to the left or right subtree of $r_{i}$. In other words, if $r_{i}$ is the ancestor of $r_{j}$, both $r_{i}\swarrow r_{j}$ and $r_{i}\searrow r_{j}$ are valid. Consequently, Axiom 4 of the proper decision tree no longer holds, and the ODT-BF problem cannot be solved using the $\mathit{odt}_{K}$ program.
As a result, ODT-BF satisfies only the structural constraints of decision trees, with no ancestral constraints. It therefore reduces to an ordinary labeled binary tree problem with no explicit constraints, yielding a complexity of $K!\times\mathit{Catalan}\left(K!\right)$ for an input of size $K$, since there are $\mathit{Catalan}\left(K!\right)$ possible tree shapes and $K!$ possible labelings.
Without ancestral constraints, the split function for ODT-BF is straightforward: once a root $r$ is fixed, how should the two sublists $\mathit{rs}^{+}$ and $\mathit{rs}^{-}$ for its subtrees be determined from $\mathit{rs}/r$? The answer is straightforward: any subset of $\mathit{rs}/r$ is valid! The following function produces all possible sublists $\mathit{xs}^{\prime}$ along with their complementary subsets $\mathit{xs}/\mathit{xs}^{\prime}$ as a pair.
$\mathit{subsPair}$ $\left(\left[\>\right],\mathit{ys}\right)=\left[\left(\left[\>\right],\mathit{ys}\right)\right]$
$\mathit{subsPair}$ $\left(x:\mathit{xs},\mathit{ys}\right)=\mathit{subsPair}\left(\mathit{xs},\mathit{ys}\right)\cup\mathit{mapL}\left(f,\mathit{subsPair}\left(\mathit{xs},\mathit{ys}\right)\right)$
where $f\left(a,\mathit{as},\mathit{bs}\right)=\left(a:\mathit{as},\mathit{bs}/a\right)$.
For instance, running $\mathit{subsPair}\left(\left[1,2,3\right],\left[1,2,3\right]\right)$ returns
$\left[\left(\left[\>\right],\left[1,2,3\right]\right),\left(\left[3\right],\left[1,2\right]\right),\left(\left[2\right],\left[1,3\right]\right),\left(\left[2,3\right],\left[1\right]\right),\left(\left[1\right],\left[2,3\right]\right),\left(\left[1,3\right],\left[2\right]\right),\left(\left[1,2\right],\left[3\right]\right),\left(\left[1,2,3\right],\left[\>\right]\right)\right]$, the $2^{3}$ possible sublists of $\left[1,2,3\right]$.
Definition 10
The split function for the ODT-BF problem is defined as
$\mathit{splits}_{\text{BF}}\left(\mathit{rs}\right)=\mathit{subsPair}\left(\mathit{rs},\mathit{rs}\right)$ (40)
Definition of the search space
With the help of splitting function, He and Little (2025b) proposed the following generator to exhaustively generate the search space for decision trees over binary feature data, as formalized below.
Definition 11
The generator for the search space of the ODT-BF problem over the binary tree datatype is defined as
$\mathit{genDTBFs}$ $:\mathcal{D}\times\left[\mathcal{R}\right]\to\left[\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)\right]$ (41)
$\mathit{genDTBFs}$ $\left(\mathit{xs},\left[\>\right]\right)=\left[\>\right]$
$\mathit{genDTBFs}$ $\left(\mathit{xs},\left[r\right]\right)=\left[\mathit{DN}\left(\mathit{DL}\left(xs^{+}\right),r,\mathit{DL}\left(xs^{-}\right)\right)\right]$
$\mathit{genDTBFs}$ $\left(\mathit{xs},\mathit{rs}\right)=\big[\mathit{DN}\left(\mathit{mapD}{}_{\cap_{\mathit{xs}^{+}}}\left(u\right),r_{i},\mathit{mapD}{}_{\cap_{\mathit{xs}^{-}}}\left(v\right)\right)\mid r_{i}\leftarrow\mathit{rs},$
$\left(\mathit{rs}^{+},\mathit{rs}^{-}\right)\leftarrow\mathit{splits}_{\text{BF}}\left(\mathit{rs}/r_{i}\right),u\leftarrow\mathit{\mathit{genDTBFs}}\left(\mathit{rs}^{+}\right),v\leftarrow\mathit{\mathit{genDTBFs}}\left(\mathit{rs}^{-}\right)\big].$
The combinatorial analysis in He and Little (2025b) showed that $\mathit{genDTBFs}$ will generate $f_{\text{BF}}\left(K\right)=K!\times\mathit{Catalan}\left(K\right)$, align with our interpretation of labeled binary tree’s complexity.
Consequently the ODT-BF problem can be defined by following brute-force program
$\mathit{odt}_{\text{BF}}\left(\mathit{xs}\right)=\mathit{min}_{E}\circ\mathit{genDTBFs}\left(\mathit{xs}\right)$ (42)
4.4.2 optimal decision tree algorithm over binary feature data
As a result, following an exactly same derivation process as the Theorem 24 by changing $\mathit{splits}$ as $\mathit{splits}_{\text{BF}}$, we can derive the following dynamic programming algorithm for solving the 42 problem.
Theorem 12
$\mathit{odt}_{\text{BF}}$ $:\mathcal{D}\times\left[\mathcal{R}\right]\to\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)$ (43)
$\mathit{odt}_{\text{BF}}$ $\left(\mathit{xs},\left[\;\right]\right)=\left[\mathit{DL}\left(\mathit{xs}\right)\right]$
$\mathit{odt}_{\text{BF}}$ $\left(\mathit{xs},\left[r\right]\right)=\left[\mathit{DN}\left(\mathit{DL}\left(xs^{+}\right),r,\mathit{DL}\left(xs^{-}\right)\right)\right]$
[…]
$r\leftarrow\mathit{rs},\left(\mathit{rs}^{+},r,\mathit{rs}^{-}\right)\leftarrow\mathit{splits}_{\text{BF}}\left(\mathit{rs}\right)\big].$
provides a solution to the brute-force specification (42).